# Engagement, User Satisfaction, and the Amplification of Divisive Content on Social Media

### v5

@@ 1.2 Ranking by stated preferences
Next, we conduct an exploratory analysis in which we simulate an alternative ranking of tweets that is based on users’ stated preferences. In particular, first, we take the approximately twenty unique tweets that each user was surveyed about and re-rank these tweets by the users’ stated preference for the tweet. Each tweet receives a score of $1$ = “Yes”, 00 = “Indifferent”, or $-1$ = “No”, depending on the users’ stated preference for the tweet. [^4] If a tweet was a quote tweet or a reply to another tweet, then we consider the two tweets to be one combined “tweet object” for the purposes of ranking and score them as the average of the users’ stated preferences for both tweets. The “Stated Preference” (SP) timeline consists of the ten tweets (out of the approximately twenty unique tweets) that scored highest according to the users’ stated preferences (ties are broken at random).
As shown in Figure 2, relative to the engagement timeline, the SP timeline reduces negativity (anger, sadness, anxiety) and increases happiness, both in terms of the emotions expressed by authors and the emotions felt by readers. Moreover, the content shown in the SP timeline is less partisan and less likely to contain expressions of out-group animosity. However, further analysis shown in Figure 2 reveals that the reduction in partisanship and animosity in the SP timeline is due almost entirely to reducing the number of tweets from the reader’s out-group, and by reducing the animosity directed toward the reader’s in-group (but not animosity towards the reader’s out-group).
@@ 2 Discussion
2.1 The effects of Twitter’s engagement-based algorithm
Amplification of emotions and out-group animosity.
Our study reveals that the engagement-based algorithm tends to amplify emotionally charged content, particularly that which expresses anger and out-group animosity. The differences were sizable. Of the political tweets chosen by Twitter’s algorithm, 62% expressed anger and 46% contained out-group animosity. In comparison, of the political tweets in the chronological timeline, 52% expressed anger and 38% contained out-group animosity. This finding helps clarify inconsistencies in previous research. Earlier studies relied on observational methods that were limited to specific subsets of content on the platform (?, ?) and replications produced mixed results (?, ?). Our randomized experiment provides insight into this debate by providing causal evidence that, indeed, the algorithm amplifies emotionally charged content. In order to ensure greater external validity, we also examined the effect of the algorithm on tweets that users themselves would actually encounter on the platform in their own timelines.
Partisan tweets, but no filter bubble.
Like prior observational and experimental work (?, ?, ?), we found that the content that the engagement-based algorithm amplified tended to be more partisan. However, we did not find evidence that the engagement-based algorithm contributed to a “filter bubble” (?). On the chronological timeline, 48% of political tweets came from the users’ in-group, and 11% came from the users’ out-group. On the other hand, on the engagement-based timeline, 50% of sociopolitical tweets were in-group tweets and 16% were out-group tweets (see Figure 2). Thus, the algorithm actually showed users more content from their out-group, and this increase was larger than the corresponding increase in in-group content. This is partially consistent with Guess et al. (2023) who found that the Facebook news feed ranking algorithm increased the amount of in-group and out-group content (?). However, they also found that the increase in in-group content was greater than the increase in out-group content—a reversal of our finding.
Amplification of polarizing content.
We asked users how they felt about their political in-group and out-group (from “Much worse” to “Much better”) after each individual piece of content they saw from either the engagement-based or chronological timeline. We found that, after reading tweets selected by the engagement-based algorithm, users tended to have more positive perceptions of their in-group and more negative perceptions of their out-group, compared to the chronological baseline. Our results suggest that the engagement-based ranking algorithm is more polarizing than a chronological one.
Nevertheless, it is unclear whether ranking algorithms have lasting effects on users’ attitudes and affective polarization. In contrast to our content-specific surveying, Guess et al. (2023) measured the effect of the Facebook and Instagram ranking algorithms on users’ affective polarization, in general, between two weeks to three months after the start of their experiment (?). They found no significant impact on users’ affective polarization. Crucially, however, the scope of Guess et al. (2023)’s research does not encompass “general equilibrium” effects, such as the possibility for ranking algorithms to indirectly shape user attitudes by incentivizing and changing the type of content that users produce in the first place.
Stated preference.
@@ 2.1 The effects of Twitter’s engagement-based algorithm
We found that users had only a slight preference for tweets in their engagement-based timeline compared to tweets in their reverse chronological timeline. This small preference is notable because Twitter’s algorithm has a large effect on how much time users spend on the platform (?, ?). Moreover, users were less likely to prefer the political tweets that the algorithms selected. This suggests that the engagement-based algorithm may be catering to users’ revealed preferences but not to their stated preferences, especially when it comes to political content.
2.2 Ranking by stated preferences
The engagement-based algorithm amplified divisive content and failed to meet users’ stated preferences. This observation led us to question: what would the implications be of ranking content by users’ stated preferences instead?
The effects of ranking by stated preference.
In exploratory analysis, we evaluated the effects of creating an alternative ranking based on users’ stated preference (SP). We found that the SP timeline significantly reduced the prevalence of angry content, partisan content, and content that expressed out-group animosity, relative to the engagement-based timeline. This suggests that part of the reason current algorithms amplify divisiveness is their prioritization of engagement over more reflective notions of preference.
@@ 2.2 Ranking by stated preferences
However, delving deeper, we also found that the decrease in partisan tweets predominantly came from a reduction in out-group tweets. Furthermore, the drop in out-group animosity was primarily due to reducing hostility towards the readers’ in-group (but not towards their out-group). This suggests that while explicitly stated preferences might lower overall animosity, they could inadvertently heighten users’ exposure to content that reinforces their pre-existing beliefs. This is in line with prior research suggesting that users’ preference for pro-attitudinal content and sources may be a primary driver of echo chambers online (?, ?, ?).
@@
Trade-offs between ranking approaches.
Given the vast amount of content available for ranking that algorithms can choose from, relative to the amount of content that an individual user can view, we hypothesize that it may be possible to do well enough with respect to all three of these desiderata: to achieve high user engagement, satisfy users’ stated preferences, and avoid amplifying hostile and polarizing content. In Section S4.9, we show that a slight preference to the stated preference timeline, one which involves tie-breaking content ranking based on the presence of out-group animosity, satisfies users’ stated preferences as much as the SP timeline, but greatly reduces the amplification of divisive content and avoids reinforcement of echo chambers. However, more research is needed to understand the trade-offs inherent between all three of these desiderata, especially with respect to engagement.
Finally, in terms of the practicalities of content ranking, it is important to note that we do not advocate for scrapping the use of engagement. Compared to user surveys, which are sparse and potentially costly to collect, engagement data is readily available to platforms and still provides some signal about what users value. Engagement will likely be an important part of any content ranking algorithm (?). But there may be smarter ways to combine dense engagement signals with sparser survey data. For example, Milli et al. (2021) created a latent variable model that learned to interpret different engagement signals in a way that remained consistent with users’ stated preferences (?).
@@ 2.3 Limitations
Our study was designed to be carefully controlled. The only difference between the two conditions in our experiment was whether the content was ranked reverse-chronologically or through the personalized algorithm. This allowed us to obtain precise estimates of the (local) effect of the algorithmic ranking. However, such an approach does not capture the algorithm’s long-term effects. For example, by incentivizing some types of content over others, the algorithm changes what type of content is produced in the first place, which, in turn, affects what is available in the chronological timeline. If users have learned to produce more of the content that the algorithm incentivizes (?), this implies that the algorithm’s long-term effects on the content-based outcomes we measure (emotions, partisanship, and out-group animosity in tweets) may be even greater. Nevertheless, rigorous experimentation needs to be done to establish how all the effects found in our study change over time in response to feedback loops (?).
Another concern pertains to the ecological validity of our study. We only surveyed participants about the first ten tweets in each timeline, a method that might not perfectly mimic real-world Twitter usage where users can scroll past various tweets. Looking at the first ten tweets might be a reasonable approximation though—Bouchaud et al. (2023) found that users viewed an average of 30 tweets in one session of using Twitter (?). In Section S4.5, we show that our results are robust to using a range of thresholds, spanning from the first five tweets to the first ten tweets.
@@
Finally, we required participants to install a Chrome extension and relied on an online crowd-working platform for participant recruitment. This may have impacted the types of participants we could attract. In particular, compared to Twitter users in the 2020 ANES study (?), our population tended to be younger (53% of our study are aged 18-34 years old, compared to 45% in the ANES study) and more likely to affiliate with the Democratic Party (56% Democrat in our study versus 41% in the ANES study) (see SM section S2 for full demographic statistics). We report on heterogenous effects by different demographic groups in Section S4.7, however, more research may be warranted before generalizing to the full Twitter population.
@@ Acknowledgements
Funding. This study was financially supported by the UC Berkeley Center for Human-Compatible AI. MC was supported by the National Science Foundation Graduate Research Fellowship Program.
@@
Competing interests. The authors declare that they have no competing interests.
Data and materials availability. The data and code needed to reproduce our results are available to the research community at http://github.com/smilli/twitter.
[…]
@@
We conducted our study between February 11 to February 27, 2023 on CloudResearch Connect, a crowd-working platform. Informed consent was obtained from all participants. The study was approved by UC Berkeley’s IRB under the CPHS protocol ID number 2021-09-14618 and complies with all ethical regulations. Moreover, at the end of the study period, our rating (given by study participants) on CloudResearch was 4.9 stars, higher than 99 percent of other researchers on the platform.
The study period was broken into four waves and participants could complete the study once during each wave. The time periods for the waves (inclusive) were 02/11-02/14, 02/16-02/19, 02/21-02/23, and 02/25-02/27. Every day, we recruited up to 150 eligible participants who lived in the United States, were at least 18 years old, and used Google Chrome. Furthermore, participants were required to use Twitter at least a few times a week and follow at least 50 people on Twitter (both gauged through self-reports). To collect data, participants were directed to download a Chrome extension that we developed which scraped their Twitter homepage to collect the top tweets from their personalized timeline. While scraping, the Chrome extension added an overlay to the homepage that prevented the user from seeing the tweets during collection. At the same time that the personalized tweets were collected, we queried the Twitter API to get the top tweets from the chronological timeline. Only public tweets were collected and no promoted tweets (advertisements) were collected.
@@
  • Political affect Left. Participants are asked: “How does [@author-handle]’s tweet make you feel about people or groups on the Left?” Participants respond on a Likert scale of “Much worse,” “Worse,” “The same as before,” “Better,”, or “Much better.”
  • Political affect Right. Participants are asked: “How does [@author-handle]’s tweet make you feel about people or groups on the Right?” Participants respond on a Likert scale of “Much worse,” “Worse,” “The same as before,”, “Better,” or “Much better,”
  • Outgroup animosity towards the left. Participants are asked: “Is [@author-handle]’s tweet expressing anger, frustration, or hostility towards a person or group on the Left?” Participants choose between a response of “Yes” or “No” This question is only asked if political leaning is “Right” or “Far right.”
  • Outgroup animosity towards the right. Participants are asked: “Is [@author-handle]’s tweet expressing anger, frustration, or hostility towards a person or group on the Right?” Participants choose between a response of “Yes” or “No.” This question is only asked if political leaning is “Left” or “Far left.”
@@
$\widebar{Y}^{\text{pol}}_{u,q}(1)=\sum_{w\in\mathcal{W}_{u}^{\text{pol}}}\frac{1}{|\mathcal{W}_{u}^{\text{pol}}|}\sum_{t\in\mathcal{T}^{\text{pol}}_{w,u}(x)}\frac{Y_{w,u,t,q}(x)}{|\mathcal{T}^{\text{pol}}_{w,u}(x)|}$ (3)
@@
Significance and multiple testing. To compute $p$-values, we conducted two-tailed paired permutation tests. A participant’s responses to their personalized tweets and responses to their chronological tweets are considered paired data, i.e., $\widebar{Y}_{u,q}(1)$ and $\widebar{Y}_{u,q}(0)$, or $\widebar{Y}^{\text{pol}}_{u,q}(1)$ and $\widebar{Y}^{\text{pol}}_{u,q}(0)$ are considered paired data. We used bootstrap sampling to calculate the 95% confidence intervals displayed in Figure 1. To adjust for the possibility of chance findings when conducting multiple tests, we used the Benjamini-Krieger-Yekutieli two-stage method (?), as described by (?), to compute false discovery rate (FDR) adjusted $p$-values. The FDR-adjusted $p$-value (sometimes called a $q$-value) for each test is the lowest false discovery rate such that the test would still be rejected at a significance level of $0.05$. We test 26 outcomes in total, and all significant results (significant at a $p$-value threshold of $0.05$) remain significant at a false discovery rate of $0.01$. Thus, on expectation, none of our discoveries are false discoveries. The full table of standardized and unstandardized effect sizes, $p$-values, and FDR-adjusted $p$-values can be found in SM section S3.
@@
In this section, we report on the demographic distribution of our participants and compare our distribution to that of Twitter users from the 2020 ANES study (?). The primary differences are that our population is younger (53 percent of our study are aged 18-34 years old, compared to 45 percent in the ANES study) and more likely to affiliate with the Democratic Party (56 percent Democrat in our study versus 41 percent in the ANES study).
@@
Users could participate in our study up to four different times, and we collected demographic information from them each time. In total, 806 unique users participated in our study 1730 times. The following tables report demographics across all users, with each user’s demographic responses being given a weight of $1/m$, where $m$ represents the number of survey waves they participated in. In addition to demographic questions, we also asked two user-level survey questions about the primary reason participants use Twitter and the primary type of content they saw in the tweets we showed them (Table 9).
Race (Our Study) Weight Percent
White
538.23 66.78
Black or African American
88.23 10.95
Hispanic
80.24 9.96
Asian or Native Hawaiian/other Pacific Islander
62.51 7.76
American Indian/Alaska Native or Other
8.84 1.10
Multiple races, non-Hispanic
27.94 3.47
@@
Gender (Our Study) Weight Percent
Man
410.26 50.90
Woman
371.00 46.03
Non-binary
20.78 2.58
Other
3.97 0.49
@@
Ideological Leaning (Our Study) Weight Percent
Far left
168.87 20.95
Left
325.56 40.39
Moderate
195.31 24.23
Right
85.86 10.65
Far right
25.29 3.14
Other
5.12 0.63
@@
Political Leaning Further (Our Study) Weight Percent
Left-leaning
126.14 62.94
Right-leaning
74.28 37.06
Total 200.42 100
@@
Summary Leaning (Our Study) Weight Percent
Left-leaning
620.57 76.99
Right-leaning
185.43 23.01
@@
Political Party (Our Study) Weight Percent
Democrat
452.51 56.14
Republican
111.08 13.78
Independent
204.42 25.36
Something else
37.99 4.71
@@
Political Party Further (Our Study) Weight Percent
Republican
69.94 28.85
Democrat
172.46 71.15
Total 242.40 100
@@
Summary Party (Our Study) Weight Percent
Republican
181.03 22.46
Democrat
624.97 77.54
@@
Education Levels (Our Study) Count Percent
Some high school
5.82 0.72
High school graduate
241.91 30.01
Associate degree
103.44 12.83
Bachelor’s degree
311.18 38.61
Master’s degree or above
127.67 15.84
Other
14.86 1.84
Prefer not to answer
1.12 0.14
@@
Age Group (Our Study) Weight Percent
18-24 years old
119.33 14.81
25-34 years old
307.24 38.12
35-44 years old
211.74 26.27
45-54 years old
92.68 11.50
55-64 years old
50.33 6.24
65-74 years old
23.67 2.94
75 years or older
1.00 0.12
@@
Annual Household Income (Our Study) Weight Percent
Less than \$25,000
117.40 14.57
\$25,000-\$50,000
189.92 23.56
\$50,000-\$100,000
301.26 37.38
\$100,000-\$200,000
147.06 18.25
More than \$200,000
34.40 4.27
Prefer not to say
15.95 1.98

### v6

@@ 2.2 Ranking by stated preferences
Next, we conducted an exploratory analysis in which we simulated an alternative ranking of tweets that is based on users’ stated preferences. In particular, we took the approximately twenty unique tweets that each user was surveyed about and re-ranked them by the users’ stated preference for the tweet. Each tweet received a score of $1$ = “Yes”, 00 = “Indifferent”, or $-1$ = “No”, depending on the users’ stated preference for the tweet. [^4] If a tweet was a quote tweet or a reply to another tweet, then we combined the two tweets into one “tweet object” and score it by the average of the users’ stated preferences for both tweets. The “Stated Preference” (SP) timeline consisted of the ten tweets (out of the approximately twenty unique tweets) that scored highest according to the users’ stated preferences (ties are broken at random). We chose the top ten tweets to maintain consistency with the engagement-based and reverse-chronological timelines (where we also considered only the top ten tweets).
As shown in Figure 1, relative to the engagement timeline, the SP timeline reduced negativity (anger, sadness, anxiety) and increased happiness, both in terms of the emotions expressed by authors and the emotions felt by readers. Moreover, the content shown in the SP timeline is less partisan and less likely to contain expressions of out-group animosity. However, further analysis shown in Figure 2 reveals that the reduction in partisanship and animosity in the SP timeline is due almost entirely to reducing the number of tweets from the reader’s out-group, and by reducing animosity directed toward the reader’s in-group (but not animosity towards the reader’s out-group).
@@ 3 Discussion
The engagement-based algorithm does not cater to users’ stated preferences.
@@
For tweets overall, we found that users had only a slight preference for tweets in their engagement-based timeline compared to tweets in their reverse chronological timeline. Moreover, users were less likely to prefer the political tweets that the engagement-based algorithm selected. This is notable because, as measured by internal A/B tests at Twitter, the engagement-based algorithm causes users to spend more time on the platform, relative to the reverse-chronological timeline [37, 7]. Thus, the engagement-based algorithm seems to be catering to users’ revealed preferences (in terms of engagement and usage patterns) but not to their stated preferences, particularly when it comes to political content. Rathje et al. [42] found that U.S. adults believe social media platforms amplify negative, emotionally-charged, and out-group hostile content but should not. This perspective may help explain why users were dissatisfied with the political tweets selected by the engagement-based algorithm, which tended to be angrier and more out-group hostile. The fact that the engagement-based algorithm is not aligned with users’ stated preference for tweets also suggests that its effects cannot entirely be explained by users’ deliberative choices.
The engagement-based algorithm amplifies partisan, emotional, out-group hostile tweets.
We found that the engagement-based algorithm tends to select more emotionally-charged, partisan, out-group hostile content than both the reverse-chronological timeline and the stated preference timeline. This finding helps clarify inconsistencies in previous research. Earlier studies relied on observational methods that were limited to specific subsets of content on the platform [11, 41] and replications produced mixed results [15, 10]. These studies found a correlation between emotion and engagement but did not control for who users followed, leaving open the possibility that the observed correlation was due to users’ choices of who to follow. Our study provides insight into this debate by providing evidence that the engagement-based algorithm amplifies emotionally charged content, beyond users’ choices of who to follow, and even beyond what users’ stated preferences at the tweet-level would indicate. It also provides evidence for the theory of Brady et al. [13] who suggest that ranking algorithms exploit humans’ natural social biases towards moralized, emotional content.
The engagement-based timeline may be more polarizing than the reverse-chronological timeline
We also asked users how they felt about their political in-group and out-group (from “Much worse” to “Much better”) after each individual piece of content they saw from either the engagement-based or reverse-chronological timeline. We found that, after reading tweets selected by the engagement-based algorithm, users tended to have more positive perceptions of their in-group and more negative perceptions of their out-group, compared to the reverse-chronological timeline. Our results suggest that the engagement-based ranking algorithm selects more polarizing content compared to what would be expected from users’ following choices alone.
However, it is unclear whether ranking algorithms have lasting effects on users’ attitudes and affective polarization. In contrast to our content-specific surveying, Guess et al. (2023) measured the effect of the Facebook and Instagram ranking algorithms on users’ affective polarization, in general, between two weeks to three months after the start of their experiment [27]. They found no significant impact on users’ affective polarization. Crucially, however, the scope of Guess et al. (2023)’s research does not encompass “general equilibrium” effects, such as the possibility for ranking algorithms to indirectly shape user attitudes by incentivizing and changing the type of content that users produce in the first place.
Ranking by stated preferences may reinforce in-group bias.
@@
We also found that while the stated preference timeline had fewer partisan tweets than the engagement-based timeline, this was primarily due to a reduction in tweets from the reader’s political out-group. Similarly, the lower out-group animosity in the stated preference timeline (relative to the engagement-based timeline) was primarily due to reducing hostility towards the readers’ in-group (but not towards their out-group). And we did not find a significant difference between the engagement-based timeline and the stated preference timeline, when it came to effects on readers’ perceptions of their in-group and out-group. Overall, these results suggests that it is possible that ranking by stated preferences could heighten users’ exposure to content that reinforces their pre-existing beliefs, relative to the engagement-based algorithm. This is in line with prior research suggesting that users’ preference for pro-attitudinal political content and sources is a key factor contributing to greater exposure to in-group content compared to out-group content [5, 25, 52].
Our findings contrast with that of Agan et al. [1]. They posit that algorithms that learn from users’ reflexive engagement tend to replicate automatic “system 1” biases. As supporting evidence, they show that, on Facebook, ranking content based on engagement increases exposure to in-group content (where in-groups are defined by race or religion), whereas ranking by stated preferences reduces this exposure. However, our results reveal an opposite trend when defining in-groups and out-groups by political affiliation. This might indicate that political in-group preferences are not solely “system 1” biases but also involve more deliberate “system 2” preferences. This distinction between political out-groups and other societal out-groups may also be an important distinction for the theory proposed by Brady et al. [13], which also posits that ranking algorithms capitalize on humans’ social biases towards in-group favoritism.
@@
It also may be that variants of the stated preference timeline can avoid reinforcing in-group bias. In Section S4.9, we investigate a slight modification to the stated preference timeline, one which involves tie-breaking content ranking based on the presence of out-group animosity, rather than random tie-breaking. This approach is similar to that of Jia et al. [33] who suggest ranking based on “societal objective functions”, e.g., an objective that prioritizes content with pro-democratic attitudes. Compared to the engagement-based and reverse-chronological timeline, we find that this modified SP timeline greatly reduces the amplification of out-group hostile content while avoiding heightened exposure to in-group content.
Implications on algorithm design.
On real-world platforms, users’ stated preferences (in the form of user surveys or user controls like a “Do not recommend” button) are rarely observed, making engagement-based signals still the predominant target for ranking (see [19] for a review of the use of non-engagement-based signals for ranking). Nonetheless, recent research in computer science has explored more effective methods for incorporating stated preference signals [36, 2, 35], making it increasingly feasible to optimize for these preferences. A fundamental, yet largely untested, assumption in this line of work is that aligning algorithms with users’ stated preferences can lead to improved outcomes for both individuals and society. Our study contributes to addressing this gap by showing that optimizing for stated preferences could decrease the prominence of emotionally charged, divisive content, which others have argued can lead to greater social misperception and conflict [13]. However, there is also a possibility that ranking by stated preferences could increase exposure to pro-attitudinal content, which warrants further investigation.
@@ 3.1 Limitations
There are several limitations to our study that should be regarded when interpreting our results. First, our study analyzed the difference in users’ engagement-based and reverse-chronological timelines during only one point in time. Such an approach does not capture the engagement-based algorithm’s long-term effects. For example, by incentivizing some types of content over others, the algorithm changes what type of content is produced in the first place, which, in turn, affects what is available in the reverse-chronological timeline. If users have learned to produce more of the content that the engagement-based algorithm incentivizes [13, 16], then its long-term effects on the content-based outcomes we measure (emotions, partisanship, and out-group animosity in tweets) may be even greater. On the other hand, other research has failed to find long-term changes in user beliefs due to social media ranking algorithms [27, 26, 38, 29], and the short-term changes in users’ perceptions of their in-group and out-group that we found may not persist. Rigorous experimentation needs to be done to establish how all the effects found in our study change over time in response to feedback loops [48].
@@
Second, we required participants to install a Chrome extension and relied on an online crowd-working platform for participant recruitment, and this may have impacted the types of participants we could attract. In particular, compared to Twitter users in the 2020 ANES study [3], our population tended to be younger (53% of our study were aged 18-34 years old, compared to 33% in the ANES study) and more likely to affiliate with the Democratic Party (56% Democrat in our study versus 43% in the ANES study) (see SM section S2 for full demographic statistics). We report on heterogenous effects by different demographic groups in Section S4.7, however, more research may be warranted before generalizing to the full Twitter population.
Another concern pertains to the ecological validity of our study. We surveyed participants about each of the first ten tweets in each timeline, a method that deviates from real-world Twitter usage where users can scroll past various tweets. Looking at the first ten tweets might be a reasonable approximation though—Bouchaud et al. (2023) found that users viewed an average of 30 tweets in one session of using Twitter [9]. In Section S4.5, we show that our results are robust to using a range of thresholds, spanning from the first five tweets to the first ten tweets. Still, the effects of real-world social media use may not be accurately captured by averaging the impact of tweets shown to users. For instance, users may only be influenced by a few particularly notable tweets as they scroll through their timeline.
There is also the possibility that respondents experienced survey fatigue, as we repeatedly asked them about twenty tweets. To address this, we included attention checks in the survey and excluded participants who failed these checks. Nonetheless, we cannot completely rule out the effect of fatigue on the study’s results. Additionally, when evaluating certain tweet-related outcomes (e.g., perceived partisanship or emotional tone) based on reader perception, participants’ own backgrounds is likely to influence their preceptions [14]. Although we conducted parallel analyses using GPT-4 to label these outcomes (see Section S4.6), large language models have their own biases [44, 20] and are limited to analyzing only the textual content of the tweets, excluding images, videos, and linked websites. The similar results from both GPT-4 and reader assessments support the robustness of our findings. Still, it is essential to consider the complementary biases from both readers and GPT-4 when interpreting both sets of results.
Finally, throughout this paper, we frequently discuss disentangling the effects of ranking algorithms from intentional human decisions such as which accounts to follow. However, over a longer period, ranking algorithms also shape these choices since users are more inclined to follow accounts that appear in their timelines. Consequently, even when using an engagement-based timeline, users’ follow decisions—shaped by the engagement-based algorithm’s recommendations—can gradually influence their reverse-chronological timeline, which is based on their followed accounts. This means that, over time, the engagement-based algorithm indirectly affects the content of the reverse-chronological timeline. Our study does not account for the potential long-term impact of the engagement-based algorithm on users’ choices of who to follow.
@@ Acknowledgements
Funding.
This study was financially supported by the UC Berkeley Center for Human-Compatible AI. MC was supported by the National Science Foundation Graduate Research Fellowship Program. Authors declare no competing interests.
@@
Data availability.
The data and code needed to reproduce our results are available to the research community at http://github.com/smilli/twitter.
@@
We conducted our study between February 11 to February 27, 2023 on CloudResearch Connect [^5] Prior academic research has found that participants recruited on CloudResearch or Prolific tend to provide higher-quality data than those recruited from Mechanical Turk or Qualtrics [21, 24]. We selected CloudResearch over Prolific because of its functionality that enables requesting that participants download a Chrome extension. CloudResearch enforces one account per participant, verifies that each IP address aligns with the reported location, and ensures that the bank or PayPal accounts for cashing out are unique to each participant. Participants are only allowed to sign up for the platform if they pass an onboarding process that checks for properties such as attention, language comprehension, honesty, and their ability to follow instructions. To maintain high data quality, CloudResearch continually monitors participants, using random attention checks and investigating those who are frequently rejected or flagged by researchers [30]., a crowd-working platform. Informed consent was obtained from all participants. The study was approved by UC Berkeley’s IRB under the CPHS protocol ID number 2021-09-14618 and complies with all ethical regulations. Moreover, at the end of the study period, our rating (given by study participants) on CloudResearch was 4.9 stars, higher than 99 percent of other researchers on the platform.
The study period was broken into four waves and participants could complete the study once during each wave. [^6] To prevent repeat participation, we ensured that each Twitter user ID could only be used once per wave. For a participant to participate multiple times in a single wave, they would have needed to create multiple Twitter accounts and bypass CloudResearch’s safeguards to register multiple CloudResearch Connect accounts. The time periods for the waves (inclusive) were 02/11-02/14, 02/16-02/19, 02/21-02/23, and 02/25-02/27. Every day, we recruited up to 150 eligible participants who lived in the United States, were at least 18 years old, and used Google Chrome. Furthermore, participants were required to use Twitter at least a few times a week and follow at least 50 people on Twitter (both gauged through self-reports). To collect data, participants were directed to download a Chrome extension that we developed which scraped their Twitter homepage to collect the top tweets from their personalized timeline. While scraping, the Chrome extension added an overlay to the homepage that prevented the user from seeing the tweets during collection. At the same time that the personalized tweets were collected, we queried the Twitter API to get the top tweets from the chronological timeline. Only public tweets were collected and no promoted tweets (advertisements) were collected.
@@
  • Political affect towards Left. Participants are asked: “How does [@author-handle]’s tweet make you feel about people or groups on the Left?” Participants respond on a Likert scale of “Much worse,” “Worse,” “The same as before,” “Better,”, or “Much better.”
  • Political affect towards Right. Participants are asked: “How does [@author-handle]’s tweet make you feel about people or groups on the Right?” Participants respond on a Likert scale of “Much worse,” “Worse,” “The same as before,”, “Better,” or “Much better,”
  • Out-group animosity towards the Left. Participants are asked: “Is [@author-handle]’s tweet expressing anger, frustration, or hostility towards a person or group on the Left?” Participants choose between a response of “Yes” or “No” This question is only asked if political leaning is “Right” or “Far right.”
  • Out-group animosity towards the Right. Participants are asked: “Is [@author-handle]’s tweet expressing anger, frustration, or hostility towards a person or group on the Right?” Participants choose between a response of “Yes” or “No.” This question is only asked if political leaning is “Left” or “Far left.”
@@
$\widebar{Y}^{\text{pol}}_{u,q}(x)=\sum_{w\in\mathcal{W}_{u}^{\text{pol}}}\frac{1}{|\mathcal{W}_{u}^{\text{pol}}|}\sum_{t\in\mathcal{T}^{\text{pol}}_{w,u}(x)}\frac{Y_{w,u,t,q}(x)}{|\mathcal{T}^{\text{pol}}_{w,u}(x)|}$ (3)
@@
Significance and multiple testing. To compute $p$-values, we conducted two-tailed paired permutation tests. A participant’s responses to their personalized tweets and responses to their chronological tweets are considered paired data, i.e., $\widebar{Y}_{u,q}(1)$ and $\widebar{Y}_{u,q}(0)$, or $\widebar{Y}^{\text{pol}}_{u,q}(1)$ and $\widebar{Y}^{\text{pol}}_{u,q}(0)$ are considered paired data. We used bootstrap sampling to calculate the 95% confidence intervals displayed in Figure 1. To adjust for the possibility of chance findings when conducting multiple tests, we used the Benjamini-Krieger-Yekutieli two-stage method [8], as described by [4], to compute false discovery rate (FDR) adjusted $p$-values. The FDR-adjusted $p$-value (sometimes called a $q$-value) for each test is the lowest false discovery rate such that the test would still be rejected at a significance level of $0.05$. We test 26 outcomes in total, and all significant results (significant at a $p$-value threshold of $0.05$) remain significant at a false discovery rate of $0.01$. Thus, on expectation, none of our discoveries are false discoveries. The full table of standardized and unstandardized effect sizes, $p$-values, and FDR-adjusted $p$-values can be found in SM section S3.
@@
In this section, we report on the demographic distribution of our participants and compare our distribution to that of Twitter users from the 2020 ANES study [3]. The primary differences are that our population is younger (53 percent of our study are aged 18-34 years old, compared to 33 percent in the ANES study) and more likely to affiliate with the Democratic Party (56 percent Democrat in our study versus 43 percent in the ANES study).
@@
Participants in our study had the opportunity to take part in up to four waves, with demographic information collected at each wave. In total, 806 unique users participated in our study 1730 times across waves. The following tables report demographics across all users. For users that participated in multiple waves, we report on their demographic information from their first wave. In addition to demographic questions, we also asked two user-level survey questions about the primary reason participants use Twitter and the primary type of content they saw in the tweets we showed them (Table 8).
Race (Our Study) $\mathbf{n}$ Percent
White 537 66.63
Black or African American 89 11.04
Hispanic 81 10.05
Asian or Native Hawaiian/other Pacific Islander 63 7.82
American Indian/Alaska Native or Other 8 0.99
Multiple races, non-Hispanic 28 3.47
@@
Gender (Our Study) $\mathbf{n}$ Percent
Man 410 50.87
Woman 371 46.03
Non-binary 21 2.61
Other 4 0.50
@@
Ideological Leaning (Our Study) $\mathbf{n}$ Percent
Far left 170 21.09
Left 322 39.95
Moderate 196 24.32
Right 86 10.67
Far right 25 3.10
Other 7 0.87
@@
Political Leaning Further (Our Study) $\mathbf{n}$ Percent
Towards the Left 136 62.67
Towards the Right 81 37.33
Total 217 100
@@
Summary Leaning (Our Study) $\mathbf{n}$ Percent
Left-leaning 618 76.67
Right-leaning 188 23.33
@@
Political Party (Our Study) $\mathbf{n}$ Percent
Democrat 451 55.96
Republican 111 13.77
Independent 205 25.43
Something else 39 4.84
@@
Political Party Further (Our Study) $\mathbf{n}$ Percent
Republican 72 29.27
Democrat 174 70.73
Total 246 100
@@
Summary Party (Our Study) $\mathbf{n}$ Percent
Republican 182 22.58
Democrat 624 77.42
@@
Education Levels (Our Study) $\mathbf{n}$ Percent
Some high school 5 0.62
High school graduate 242 30.02
Associate degree 102 12.66
Bachelor’s degree 310 38.46
Master’s degree or above 128 15.88
Other 17 2.11
Prefer not to answer 2 0.25
@@
Age Group (Our Study) $\mathbf{n}$ Percent
18-24 years old 120 14.89
25-34 years old 307 38.09
35-44 years old 211 26.18
45-54 years old 93 11.54
55-64 years old 51 6.33
65-74 years old 23 2.85
75 years or older 1 0.12
@@
Annual Household Income (Our Study) $\mathbf{n}$ Percent
Less than \$25,000 118 14.64
\$25,000-\$50,000 188 23.33
\$50,000-\$100,000 300 37.22
\$100,000-\$200,000 149 18.49
More than \$200,000 35 4.34
Prefer not to say 16 1.99